### v1

@@ 5.3 Computational experiments with synthetic datasets
2
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.00 (4.49)
94.49 (3.02)
92.87 (3.38)
99.94 (0.16)
100 (0.00)
84.62 (5.23)
89.57 (4.04)
88.51 (3.99)
97.97 (1.01)
100 (0.00)
2.92 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.53 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.98 (0.07)
2.17 (0.21)
2.0 (0.00)
1.71 (0.25)
1.70 (0.25)
4
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
79.94 (4.53)
84.17 (4.41)
82.80 (4.16)
95.13 (1.38)
100 (0.00)
73.97 (4.74)
76.72 (4.48)
76.16 (4.82)
90.97 (2.55)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.10 (0.3)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.48 (0.13)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.24 (0.18)
2.00 (0.00)
1.67 (0.13)
1.67 (0.12)
8
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
75.05 (5.72)
79.25 (5.66)
77.78 (5.36)
89.08 (4.26)
100 (0.00)
67.54 (7.08)
69.66 (6.75)
69.40 (6.37)
82.95 (5.70)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.12 (0.325)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.48 (0.14)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.21 (0.22)
2.00 (0.00)
1.48 (0.12)
1.49 (0.13)
12
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
70.39 (3.82)
74.50 (4.00)
73.35 (3.49)
85.06 (5.19)
100 (0.00)
62.78 (5.20)
64.03 (4.99)
63.53 (4.78)
76.71 (6.47)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.14 (0.35)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.46 (0.14)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.26 (0.13)
2.00 (0.00)
1.50 (0.11)
1.52 (0.10)
Figure 10: Effects of dimensionality. Training size = 100. No noise. Ground truth tress are size 2.
@@
Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
@@
0
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.42 (4.61)
95.19 (3.11)
93.49 (3.60)
99.62 (0.68)
1 (0.00)
87.25 (5.20)
91.21 (3.94)
89.99 (3.71)
98.06 (1.21)
1 (0.00)
2.92 (0.27)
4.0 (0.00)
2.98 (0.04)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.49 (0.16)
1.99 (0.05)
1.67 (0.00)
1.67 (0.00)
1.99 (0.03)
2.11 (0.23)
1.98 (0.11)
1.67 (0.28)
1.68 (0.28)
5
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
86.29 (4.48)
90.64 (3.07)
89.07 (3.45)
95.00 (0.20)
0.95 (0.00)
87.15 (4.89)
91.01 (3.67)
89.87 (3.47)
97.20 (1.40)
1 (0.00)
2.94 (0.24)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.2.00 (0.40)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.98 (0.08)
2.50 (0.16)
2.00 (0.00)
1.68 (0.00)
1.67 (0.00)
1.99 (0.02)
2.10 (0.24)
2.00 (0.00)
1.62 (0.29)
1.68 (0.28)
10
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
82.25 (3.78)
86.25 (2.67)
84.69 (3.09)
90.45 (0.49)
0.90 (0.00)
86.74 (5.36)
90.45 (4.07)
89.48 (3.96)
97.59 (1.21)
100 (0.00)
2.98 (0.14)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.12 (0.32)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
3.12 (0.32)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.16 (0.26)
2.00 (0.00)
1.64 (0.30)
1.68 (0.28)
15
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
77.90 (3.99)
81.80 (3.10)
80.53 (3.04)
85.69 (0.54)
84.99 (0.00)
85.99 (5.71)
90.19 (3.67)
89.10 (3.79)
97.10 (1.84)
100 (0.00)
2.98 (0.14)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.22 (0.41)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.51 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.04)
2.19 (0.31)
2.00 (0.00)
1.63 (0.29)
1.68 (0.28)
20
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
73.88 (3.16)
77.36 (2.91)
76.28 (2.81)
80.98 (0.69)
79.90 (0.00)
86.06 (4.95)
89.51 (4.59)
88.61 (4.43)
96.58 (2.06)
100 (0.00)
2.94 (0.24)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.98 (0.08)
2.54 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.13 (0.34)
2.00 (0.00)
1.68 (0.27)
1.68 (0.28)
25
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
70.35 (3.22)
73.19 (2.54)
72.26 (2.62)
76.37 (0.83)
75.00 (0.00)
84.59 (6.66)
87.78 (5.41)
87.06 (5.16)
95.67 (2.93)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.52 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.24 (0.32)
2.00 (0.00)
1.67 (0.78)
1.68 (0.28)
@@
As the noise level increased, the accuracy of all methods tended to decrease, with similar effects on out-of-sample performance up to a noise level of 20%. Beyond this point (from 20% to 25% noise), differences became more pronounced: CART-depth, CART-size, and ConTree decreased by 1.47%, 1.73%, and 1.55%, respectively, whereas HODT decreased by only 0.91%. From a broader perspective, when increasing noise from 0% to 25%, HODT’s out-of-sample accuracy decreased by 2.39%, compared with 2.93%, 3.43%, and 2.66% for ConTree, CART-size, and CART-depth, respectively. Notably, starting at a 10% noise level, HODT was able to find solutions that exceeded the ground truth on the training dataset while maintaining strong performance on out-of-sample tests. These results further refute the notion that optimal methods are less robust to noise, highlighting the superiority of HODT.
@@
Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
@@
0
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.29 (4.74)
94.20 (3.70)
92.73 (3.90)
99.88 (0.26)
100 (0.00)
87.23 (5.14)
90.32 (4.24)
89.55 (4.10)
98.08 (1.00)
100 (0.00)
2.96 (0.20)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.2 (0.399)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.065)
2.51 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.09 (0.33)
2.00 (0.00)
1.63 (0.29)
1.63 (0.28)
5
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.16 (4.63)
93.98 (3.78)
92.62 (3.79)
99.62 (0.40)
99.59 (0.44)
87.16 (5.06)
90.32 (4.20)
89.50 (3.94)
98.10 (1.01)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.53 (0.168)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.05 (0.32)
2.00 (0.00)
1.64 (0.28)
1.65 (0.28)
10
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.30 (4.55)
93.98 (3.57)
92.50 (3.65)
99.07 (0.62)
98.81 (0.71)
87.36 (5.09)
90.30 (4.55)
89.38 (4.09)
97.50 (0.90)
100 (0.00)
2.96 (0.20)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.28 (0.45)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.065)
2.54 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.03 (0.34)
2.00 (0.00)
1.58 (0.29)
1.65 (0.28)
15
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.86 (4.91)
93.78 (3.92)
92.52 (3.83)
99.05 (0.66)
98.72 (0.80)
87.06 (4.97)
90.27 (4.42)
89.31 (4.38)
97.90 (1.10)
100 (0.00)
2.96 (0.20)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.32 (0.47)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.07)
2.55 (0.18)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.02)
2.03 (0.33)
2.00 (0.00)
1.63 (0.28)
1.65 (0.28)
20
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.75 (4.98)
93.75 (3.70)
92.39 (3.72)
98.65 (0.87)
98.11 (1.04)
86.61 (5.57)
90.21 (4.27)
89.25 (4.05)
97.40 (11.50)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.49 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.08 (0.27)
2.00 (0.00)
1.64 (0.28)
1.65 (0.28)
25
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.76 (4.47)
93.58 (4.13)
92.29 (3.65)
98.35 (1.00)
97.67 (0.91)
87.13 (4.93)
90.28 (4.50)
89.55 (4.17)
97.62 (1.20)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
2.98 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.046)
2.53 (0.17)
1.99 (0.05)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.05 (0.35)
1.99 (0.05)
1.67 (0.27)
1.65 (0.28)
Figure 11: The effect of noise on training data. Training data size = 100. Ground truth trees are size 2.

### v2

@@ 5.3 Computational experiments with synthetic datasets
2 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.00 (4.49) 94.49 (3.02) 92.87 (3.38) 99.94 (0.16) 100 (0.00) 84.62 (5.23) 89.57 (4.04) 88.51 (3.99) 97.97 (1.01) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.98 (0.07) 2.17 (0.21) 2.0 (0.00) 1.71 (0.25) 1.70 (0.25)
4 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 79.94 (4.53) 84.17 (4.41) 82.80 (4.16) 95.13 (1.38) 100 (0.00) 73.97 (4.74) 76.72 (4.48) 76.16 (4.82) 90.97 (2.55) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.10 (0.3) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.48 (0.13) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.24 (0.18) 2.00 (0.00) 1.67 (0.13) 1.67 (0.12)
8 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 75.05 (5.72) 79.25 (5.66) 77.78 (5.36) 89.08 (4.26) 100 (0.00) 67.54 (7.08) 69.66 (6.75) 69.40 (6.37) 82.95 (5.70) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.12 (0.325) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.48 (0.14) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.21 (0.22) 2.00 (0.00) 1.48 (0.12) 1.49 (0.13)
12 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 70.39 (3.82) 74.50 (4.00) 73.35 (3.49) 85.06 (5.19) 100 (0.00) 62.78 (5.20) 64.03 (4.99) 63.53 (4.78) 76.71 (6.47) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.14 (0.35) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.46 (0.14) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.26 (0.13) 2.00 (0.00) 1.50 (0.11) 1.52 (0.10)
Table 3: Effects of dimensionality. Training size = 100. No noise. Ground truth tress are size 2.
@@
Noise level Method Train Acc (%) Test Acc (%) Tree size Depth
@@
0 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.42 (4.61) 95.19 (3.11) 93.49 (3.60) 99.62 (0.68) 100 (0.00) 87.25 (5.20) 91.21 (3.94) 89.99 (3.71) 98.06 (1.21) 100 (0.00) 2.92 (0.27) 4.0 (0.00) 2.98 (0.04) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.49 (0.16) 1.99 (0.05) 1.67 (0.00) 1.67 (0.00) 1.99 (0.03) 2.11 (0.23) 1.98 (0.11) 1.67 (0.28) 1.68 (0.28)
5 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 86.29 (4.48) 90.64 (3.07) 89.07 (3.45) 95.00 (0.20) 95.00 (0.00) 87.15 (4.89) 91.01 (3.67) 89.87 (3.47) 97.20 (1.40) 100 (0.00) 2.94 (0.24) 4.0 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.2.00 (0.40) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.98 (0.08) 2.50 (0.16) 2.00 (0.00) 1.68 (0.00) 1.67 (0.00) 1.99 (0.02) 2.10 (0.24) 2.00 (0.00) 1.62 (0.29) 1.68 (0.28)
10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 82.25 (3.78) 86.25 (2.67) 84.69 (3.09) 90.45 (0.49) 90.00 (0.00) 86.74 (5.36) 90.45 (4.07) 89.48 (3.96) 97.59 (1.21) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.12 (0.32) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 3.12 (0.32) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.16 (0.26) 2.00 (0.00) 1.64 (0.30) 1.68 (0.28)
15 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 77.90 (3.99) 81.80 (3.10) 80.53 (3.04) 85.69 (0.54) 84.99 (0.00) 85.99 (5.71) 90.19 (3.67) 89.10 (3.79) 97.10 (1.84) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.22 (0.41) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.51 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.04) 2.19 (0.31) 2.00 (0.00) 1.63 (0.29) 1.68 (0.28)
20 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 73.88 (3.16) 77.36 (2.91) 76.28 (2.81) 80.98 (0.69) 79.90 (0.00) 86.06 (4.95) 89.51 (4.59) 88.61 (4.43) 96.58 (2.06) 100 (0.00) 2.94 (0.24) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.98 (0.08) 2.54 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.13 (0.34) 2.00 (0.00) 1.68 (0.27) 1.68 (0.28)
25 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 70.35 (3.22) 73.19 (2.54) 72.26 (2.62) 76.37 (0.83) 75.00 (0.00) 84.59 (6.66) 87.78 (5.41) 87.06 (5.16) 95.67 (2.93) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.52 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.24 (0.32) 2.00 (0.00) 1.67 (0.78) 1.68 (0.28)
@@
As the noise level increased, the accuracy of all methods tended to decrease, with similar effects on out-of-sample performance up to a noise level of 20%. Beyond this point (from 20% to 25% noise), differences became more pronounced: CART-depth, CART-size, and ConTree decreased by 1.47%, 1.73%, and 1.55%, respectively, whereas HODT decreased by only 0.91%. From a broader perspective, when increasing noise from 0% to 25%, HODT’s out-of-sample accuracy decreased by 2.39%, compared with 2.93%, 3.43%, and 2.66% for ConTree, CART-size, and CART-depth, respectively. Notably, starting at a 5% noise level, HODT was able to find solutions that exceeded the ground truth on the training dataset while maintaining strong performance on out-of-sample tests. These results further refute the notion that optimal methods are less robust to noise, highlighting the superiority of HODT.
@@
Noise level Method Train Acc (%) Test Acc (%) Tree size Depth
@@
0 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.29 (4.74) 94.20 (3.70) 92.73 (3.90) 99.88 (0.26) 100 (0.00) 87.23 (5.14) 90.32 (4.24) 89.55 (4.10) 98.08 (1.00) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.2 (0.399) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.065) 2.51 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.09 (0.33) 2.00 (0.00) 1.63 (0.29) 1.63 (0.28)
5 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.16 (4.63) 93.98 (3.78) 92.62 (3.79) 99.62 (0.40) 99.59 (0.44) 87.16 (5.06) 90.32 (4.20) 89.50 (3.94) 98.10 (1.01) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.05 (0.32) 2.00 (0.00) 1.64 (0.28) 1.65 (0.28)
10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.30 (4.55) 93.98 (3.57) 92.50 (3.65) 99.07 (0.62) 98.81 (0.71) 87.36 (5.09) 90.30 (4.55) 89.38 (4.09) 97.50 (0.90) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.28 (0.45) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.065) 2.54 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.03 (0.34) 2.00 (0.00) 1.58 (0.29) 1.65 (0.28)
10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.86 (4.91) 93.78 (3.92) 92.52 (3.83) 99.05 (0.66) 98.72 (0.80) 87.06 (4.97) 90.27 (4.42) 89.31 (4.38) 97.90 (1.10) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.32 (0.47) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.07) 2.55 (0.18) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.02) 2.03 (0.33) 2.00 (0.00) 1.63 (0.28) 1.65 (0.28)
20 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.75 (4.98) 93.75 (3.70) 92.39 (3.72) 98.65 (0.87) 98.11 (1.04) 86.61 (5.57) 90.21 (4.27) 89.25 (4.05) 97.40 (11.50) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.49 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.08 (0.27) 2.00 (0.00) 1.64 (0.28) 1.65 (0.28)
25 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.76 (4.47) 93.58 (4.13) 92.29 (3.65) 98.35 (1.00) 97.67 (0.91) 87.13 (4.93) 90.28 (4.50) 89.55 (4.17) 97.62 (1.20) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 2.98 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.046) 2.53 (0.17) 1.99 (0.05) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.05 (0.35) 1.99 (0.05) 1.67 (0.27) 1.65 (0.28)
Table 5: The effect of noise on training data. Training data size = 100. Ground truth trees are size 2.